Consider two polluting firms that want to abate 100 units of pollution together, in the cheapest way possible. Their total abatement costs are given below:

TAC1 (x1) = x12    & TAC2 (x2) = 3x22

  1. (1 point) Find an equation for the Aggregate Total Abatement Cost that combines the two firms (ATAC). In other words, sum TAC1 and TAC2.
  2. (1 point) Using the equality x1 + x2 = 100, find the ATAC equation in terms of x1 only.
  3. (1 point) What values of x1 and x2 minimize ATAC? Hint: find x1 first.
  4. (1 point) Calculate the Marginal Abatement Costs for each firm, at the level of x1 and x2 you found in the previous question. How does MAC1 relate to MAC2?
Answer #1

a)

Aggregate total abatement cost will be the sum of invidual firms' total abatement costs. So,

ATAC=TAC1+TAC2

ATAC=x_{1}^{2}+3x_{2}^{2}

b)

We are given that

11 + 12 = 100

On rearranging, we get

x_{2}=100-x_{1}

So,

AT AC = x + 3(100 – 21)

ATAC=x_{1}^{2}+3(10000-200x_{1}+x_{1}^{2})

ATAC=4x_{1}^{2}-600x_{1}+30000

c)

To minimize ATAC, differentiate ATAC with respect to x1 and set is equal to 0

DAT AC/dri = 8.01 – 600

Set dTATC/dx1=0

8x_{1}-600=0

x1=75 (ATAC minimizing value of x1)

x2=100-x1=100-75=25 (ATAC minimizing value of x1)

d)

MAC_{1}=dATAC_{1}/dx_{1}=2x_{1}

Set x1=75

MAC1=2*75=$150

Now

MAC_{2}=dATAC_{2}/dx_{2}=6x_{2}

Set x2=25

MAC2=6*25=$150

We find that MAC1=MAC2 at level of x1 and x2 found in earlier part.